OpenStudy (anonymous):

The length of a rectangle is 2m more than its width. If the length is increased by 6 m and the width is decreased by 2m , the area becomes 119m. Find the dimensions of the rectangle

OpenStudy (anonymous):

this seems easy but i don't know why i m not getting it

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Suppose length is \(a\) and width is \(b\) The length of a rectangle is 2m more than its width,so: \(a =b+ 2\) (1) And If the length is increased by 6 m and the width is decreased by 2m , the area becomes 119m: \((a+6)\times(b-2)=119\) \(\Leftrightarrow 2a - 6b -ab = -131\) (2) From (1),(2) \(\Rightarrow \begin{cases} b=9m \\ a=11m \end{cases}\)

OpenStudy (anonymous):

thank you!